### v1

# LaB-RAG: Label Boosted Retrieval Augmented Generation for Radiology Report Generation

Abstract
In the current paradigm of image captioning, deep learning models are trained to generate text from image embeddings of latent features. We challenge the assumption that these latent features ought to be high-dimensional vectors which require model fine tuning to handle. Here we propose Label Boosted Retrieval Augmented Generation (LaB-RAG), a text-based approach to image captioning that leverages image descriptors in the form of categorical labels to boost standard retrieval augmented generation (RAG) with pretrained large language models (LLMs). We study our method in the context of radiology report generation (RRG), where the task is to generate a clinician’s report detailing their observations from a set of radiological images, such as X-rays. We argue that simple linear classifiers over extracted image embeddings can effectively transform X-rays into text-space as radiology-specific labels. In combination with standard RAG, we show that these derived text labels can be used with general-domain LLMs to generate radiology reports. Without ever training our generative language model or image feature encoder models, and without ever directly “showing” the LLM an X-ray, we demonstrate that LaB-RAG achieves better results across natural language and radiology language metrics compared with other retrieval-based RRG methods, while attaining competitive results compared to other fine-tuned vision-language RRG models. We further present results of our experiments with various components of LaB-RAG to better understand our method. Finally, we critique the use of a popular RRG metric, arguing it is possible to artificially inflate its results without true data-leakage. Our code is available at: https://github.com/uc-cdis/label-boosted-RAG-for-RRG.
11footnotetext: These authors contributed equally22footnotetext: Corresponding author: rgrossman1@uchicago.edu11footnotetext: Department of Computer Science, University of Chicago22footnotetext: Medical Scientist Training Program, Pritzker School of Medicine33footnotetext: Center for Translational Data Science, University of Chicago44footnotetext: Department of Medicine, University of Chicago00footnotetext: Preprint. Under Review.
1 Introduction
Figure 1: Overview of LaB-RAG for RRG.
Radiology reports are free-text natural language notes describing the observations seen in radiological images, such as X-rays, CT scans, or MRI scans. These reports are written by board-certified radiologists, doctors who have had years of medical training to read these imaging modalities. This is a critically important task as radiological imaging provides a noninvasive view of internal organs and tissues, offering clinicians insights into how best to intervene and treat disease. However, the skills required to read these images can be highly specialized [67]; there is also an increasing demand for radiologists to read the growing number of images collected [39, 13, 66]. While remote radiology increased with the COVID-19 pandemic, these services still face challenges affecting the limited radiologist workforce [26]. Motivated by these considerations and coupled with the popularization of large language models (LLMs), there has been an increasing interest in developing AI tools to help bridge the gap for the radiologist shortage [24, 52, 38].
Radiology report generation (RRG) is the task of automatically generating these reports given the image [69, 47]. While RRG can be applied in any radiological imaging modality, the most frequent modality studied in the literature is X-rays of the chest. This is likely due to the frequent and inexpensive collection of chest X-rays (CXRs) in the clinical setting and the public availability of paired chest X-ray and report datasets [69], such as MIMIC-CXR [35, 36], OpenI [14], PadChest [7], BIMCV-COVID19 [77], VinDr-CXR [55], and others. For CXRs, the task is typically formulated as generating the “Findings”, “Impression”, or both sections of the corresponding report [47]. Conceptually, the findings section describes all positive or negative observations seen in the X-ray, while the impression section summarizes and interprets those findings with recommendations for clinical diagnosis [59].
At its core, RRG is a form of image captioning, specialized for medical imaging [71]. In recent years, LLMs have demonstrated impressive performance across the medical domain, passing and excelling at US medical board exams and outperforming clinicians in certain tasks [58, 11, 68]. While these results do not translate directly to clinical viability, medical image captioning is a natural area of research application for such LLMs. Efforts typically involve fine-tuning vision-language models for the medical image modality of interest [3, 1]. Even when starting with a pre-trained language, vision, or multimodal foundation models (FMs), domain adaptation is typically expected when the task is specific and potentially out-of-distribution [6].
Model adaptation classically involves model training via supervised fine tuning (SFT). SFT of FMs is becoming increasingly difficult as models become larger, requiring compute resources greater than consumer-grade workstations can provide [75]. While parameter-efficient fine tuning (PEFT) methods have demonstrated competitive results [16], a form of LLM adaptation that does not require model training is in-context learning (ICL) [17]. The goal of ICL is to have an LLM infer the target output using examples of input-output pairs given jointly at inference time with the target input [49]. Related to ICL is retrieval augmented generation (RAG), a framework for providing additional context to an LLM prompt by retrieving documents related to the input query [42]. ICL and RAG can be combined to retrieve examples that are specific to the target input [19]. However, the standard approach for ICL and RAG uses text examples and inputs, whereas RRG is an image to text task.
While general domain vision-language FMs are improving even on medical vision-language tasks [68, 48], there are an increasing number of medical modality specific FMs which may offer better performance for tasks over these modalities [51, 78, 80, 23, 74, 53, 10, 46, 4, 2, 21, 5, 81]. It is an open area of research on how to compose together these FMs which were not necessarily jointly trained [9, 45]. Furthermore, the composition or adaptation method depends on the task. For image captioning, text generation must generally be informed by the image features. We argue that the image features need not be high-dimensional latent or token embeddings, as in the current paradigm of multimodal LLMs.
Comparison LaB- RAG RAG SFT
SOTA on clinical metrics
No fine tuning of DL models
Uses disjoint vision/text models
Modular inference components
Simple model ensemble
Table 1: Conceptual comparison of LaB-RAG with other standard frameworks.
We propose LaB-RAG, Label Boosted Retrieval Augmented Generation, a method for image captioning via RAG and ICL without the need to train any deep learning models. We study LaB-RAG in the context of RRG. Fig. 1 shows a visual overview of LaB-RAG compared to standard RAG. Tab. 1 includes a high-level conceptual comparison of LaB-RAG to standard RAG and SFT methods for RRG.
Main Contributions:
  • A generalizable framework to use classical machine learning (ML) to boost LLM performance. By training simple ML models, e.g. logistic regression, to derive categorical labels from images, we propose to use the labels to inform RAG retrieval and to use the label names as textual image descriptors input to LLMs. This adapts from frozen pretrained vision models to frozen pretrained LLMs, only ever training classical models.
  • State of the art (SOTA) performance for RRG. We show that LaB-RAG for RRG can achieve SOTA when compared with RRG models from the literature. We further experiment with and analyze effects of different components of LaB-RAG.
  • Critique of a commonly used RRG metric. We critique the use of a label-based RRG metric commonly used in the literature, demonstrating that we can inflate its results by using its labels in LaB-RAG.
2 Related Work
Figure 2: LaB-RAG inference for RRG. Symbols correspond to those in Alg. 1.
Algorithm 1 Pseudocode of LaB-RAG for RRG
[…]
2:retrieval studies with image-label-text tuples
$S\leftarrow\{(i_{x},l_{x},t_{x})\}$
3:image embedding model $\theta_{img}$
4:text generative model $\phi_{txt}$
[…]
6:compute retrieval embeddings
$E\leftarrow[...,\vec{e}_{x},...]^{\intercal}$,  $\vec{e}_{x}\leftarrow\theta_{img}(i_{x})$
7:binarize retrieval labels $B\leftarrow\{b_{x}\leftarrow l_{x}=1\}$
8:train image classification model $\psi_{img}(E)\rightarrow B$
[…]
10:compute image cosine similarity
[…]
11:sort studies by similarity $\vec{r}\leftarrow[...,s_{x},...]$,  $d_{x}\geq d_{x+1}$
12:Label Filter: filter or rerank $s_{x}\in\vec{r}$ by comparing
[…]
13:retrieve studies $S_{k}$ of the $k$ highest ranked samples in $\vec{r}$
[…]
As interest in RRG has steadily increased [69], there are an abundance of available RRG models from the literature trained to generate reports over CXRs. Additionally, the recent BioNLP workshop at ACL 2024 hosted a shared task on RRG [15] where several new models were presented. These published methods can be categorized by the report sections they generate, the “Findings” [73], “Impression” [18, 64, 28, 54, 65], both independently [12, 56], or both jointly [72]. Models from the literature can be further divided by the method for generation, either by a generative LLM conditioned only on image features [56, 12, 73, 54] or by text retrieval and processing [18, 64, 28, 54, 72, 65].
There are two primary ways by which LLMs are fine-tuned to generate the report based on input CXR embeddings. CXRMate [56] is trained with image embeddings input via cross attention [76, 8, 44]. CheXagent [12] and RGRG [73] are instead trained with image embeddings prepended as input tokens before the report, adapting the image embeddings into token embedding space.
The following retrieval-based methods use cross-modal image-to-text retrieval, requiring training of a retrieval model with a joint embedding space for CXRs and their corresponding reports. CXR-RePaiR-Gen [65] leverages CXR-ReDonE [64] for its cross-modal retrieval models and otherwise is the closest implementation of a standard RAG pipeline. FactMM-RAG [72] also employs RAG for inference, but trains its own retrieval model using RadGraph [27] labels to inform the joint embedding space. This is broadly similar to the training of X-REM’s [28] retrieval model, however X-REM uses CheXbert [70] labels. X-REM outputs a concatenation of retrieved text as the final report. CXR-RePaiR [18] also uses concatenation of retrieved text for its final output, however its retrieval model is trained via the basic CLIP [62] method. CXR-ReDonE [64] is the same as CXR-RePaiR except the training/retrieval data was cleaned to remove “priors” indicating a previous CXR [63].
The most closely related method compared to LaB-RAG is Pragmatic Retrieval/Llama [54]. Like LaB-RAG, Pragmatic derives categorical labels directly from the CXR. However, Pragmatic trains an end-to-end ResNet50 [22] model, whereas LaB-RAG uses simpler logistic classifiers trained over extracted image embeddings. Additionally, Pragmatic requires the report’s “Indication”, the clinical reason and motivation for the imaging study. While LaB-RAG uses both image embedding similarity and label matching for retrieval, Pragmatic Retrieval only uses label matching of image and indication for retrieval of report text. Pragmatic Retrieval then concatenates label-retrieved text as the final report. Pragmatic Llama does no retrieval, instead training an LLM to generate the report given the indication text and the positive image labels as text. LaB-RAG also uses labels as textual image descriptors but relies on ICL and model instructing, and thus does not require model fine-tuning.
3 LaB-RAG Framework
Figure 3: LaB-RAG does better than other retrieval methods (CXR-RePaiR/ReDonE, X-REM) and no different than SFT methods on “Impression” generation (CheXagent, CXRMate) and comparable to “Findings” generation measured by F1CheXbert (RGRG, CheXagent, CXRMate). See Sec. 5.3 for analysis of F1RadGraph results for “Findings” generation.
LaB-RAG uses a label boosted RAG algorithm with ICL to do image captioning. LaB-RAG retrieves paired example text using image embedding similarity. Retrieved texts are then fed into a general domain LLM with strong instruction following and natural language comprehension capabilities. We improve upon standard RAG by incorporating predicted categorical labels into both the example retrieval and text augmentation steps. The overview of LaB-RAG applied to RRG is presented in Fig. 2 with its high-level pseudocode described in Alg. 1. We evaluate our generation results against ground truth radiology reports using natural language metrics and clinical radiology language metrics. We compare our method against models from literature and against variations of our method.
3.1 Image, Text, and Label Dataset
For our study, we use chest X-rays, radiology reports, CheXpert [25] labels, metadata, and data splits from MIMIC-CXR v2.1.0 and its related datasets. [36, 32, 35, 33]. The dataset is split at the patient level. The final number of samples used in each of our experiments depends on the availability of all required data modalities. Further details on the dataset are provided in Sec. S2.
Because the radiology reports are written at the study level and the CheXpert [25] labeler derives labels from the report, the categorical labels are also defined per study. In order of preference, CheXpert labels are extracted from the “Impression”, “Findings”, or last paragraph sections of the report [32]. CheXpert provides 14 labels where each label describes a radiological observation, including “No Finding”. Each label gets a value of 1 (positive), 0 (negative), -1 (uncertain), or null (unmentioned). A report receives a multilabel classification where any combination of labels and values can be assigned. It is thus possible to have a study with no positive labels; in such cases, we assign a positive “Other” label which is negative in all other instances. Our final labels are sets of 15 labels per study.
3.2 Image Feature Extraction
We compute image feature embeddings to train our image classifier and to enable retrieval of visually similar images and their associated text. For CXR data, we use the pretrained X-ray image encoder of BioViL-T [2] to extract image features. We select BioViL-T as our image encoder because it was trained over MIMIC-CXR, it achieved impressive performance on tasks via linear probing, and because the hi-ml-multimodal v0.2.2 python package made it simple to integrate the model into our code. While the joint vision-language training may have helped BioViL-T learn more meaningful embeddings for downstream tasks, LaB-RAG is agnostic to the method by which the image encoder was trained. Specifically, we extract X-ray embeddings with the frozen image-encoder of BioViL-T. For image-based retrieval, we use projected 128d embeddings; for training our image classifier, we use unprojected 512d embeddings.
3.3 Training LaB-Classifiers for Label Prediction
Because the CheXpert [25] labels are computed using the radiology report, using these labels of the target X-ray in the generation of its corresponding report constitutes a form of data-leakage. We train a set of logistic regression models, LaB-Classifiers, on frozen image embeddings to classify images directly, thereby preventing this leakage (Fig. S1). Further LaB-Classifier training details are in Sec. S3.
3.4 Label Boosted RAG Algorithm
To generate captions from images, LaB-RAG uses RAG with retrieved example text for ICL. We enhance both retrieval and augmentation steps using categorical labels describing the images and their corresponding text. Given an image at inference time, we rank the similarity of the inference image with all retrieval images in image embedding space (Fig. 2 Top). We apply label-based logic to filter or rerank the similarity scores (Fig. 2 Middle), described in Sec. 3.4.1. We retrieve the corresponding text of the highest ranked images, augment a prompt with the retrieved examples, and autoregressively generate a caption using a pretrained LLM (Fig. 2 Bottom). We additionally incorporate labels in the prompt formatting step (Fig. 2 Middle), described in Sec. 3.4.2. See Alg. 1 for pseudocode.
For RRG, LaB-RAG by default uses BioViL-T [2] projected image embeddings, an “Exact” label filter, the top-5 ranked examples, the “Simple” label format and prompt, and Mistral-7B-Instruct-v0.3 [50]. For generation, we serve the LLM using vLLM [40] and do greedy decoding up to 512 tokens . In all experiments, we use the training and validation splits of our data as our retrieval set and do inference over the test split. Given a retrieval corpus of the target report section, LaB-RAG is able to generate any arbitrary section; however, depending on the radiology report section(s) we aim to generate and compare with, we filter the retrieval and inference sets to only studies with the target section.
3.4.1 Label Boosted Filtering
LaB-RAG does binary label matching to filter or rerank the ranked list of retrieval samples. LaB-RAG’s label boosting module takes an input list of samples, ranked by image similarity to the inference image, and outputs a ranked list of samples (Alg. 1 Step 8). We present three variations of this module. The simplest variant is “No-filter”, where we do not perform label-based filtering or reordering.
“Exact” filtering requires that every retrieved sample has the same binary labels as the inference image’s labels:
[…]
“Partial” filtering relaxes the constraint of the exact filter by re-sorting the retrieved samples based on the amount of overlapping positive labels between each sample’s label and the inferred image label:
[…]
3.4.2 Label Boosted Formatting and Prompts
LaB-RAG uses categorical label names as textual image descriptors for image captioning. It does so by formatting the labels as text in the prompt for both the text examples and the inference image. Because the label formatting is intricately associated with the prompt structure, we include the description of our prompts here. In the context of RRG, each CheXpert [25] label is a 14 multiclass multilabel with a 15th binary class label of exclusion (our “Other” label). We thus present four label format and prompt combinations varying the degree of label verbosity and label type description and instruction. The “Naive” prompt does not utilize label descriptors and is closest to a standard RAG prompt with some model instructions.
The “Simple” prompt formats positive labels as a comma separated list before each example of text. It additionally applies the label formatting to the predicted labels of the inference image and appends this label text to the end of the prompt to condition the generation of the image’s report; the instructions within the prompt include minimal further details describing the label.
The “Verbose” prompt additionally includes all label values (negative, uncertain, and unmentioned for CheXpert labels) and expands the model instructions to describe these value types. The prompt does not describe the labels themselves. The “Instruct” prompt uses the exact same template as “Verbose” but adds explicit instructions on how to handle each value type. As the predicted image labels are binary positive or negative, the label set for the inference image will never have labels listed under other values. We refer the reader to Sec. S5 for the examples of the exact prompts and label formats.
Figure 4: Comparison of simple individual label boosting components of LaB-RAG. Compared to SFT methods (Fig. 3), we attain a comparable ${\sim}5\%$ improvement in F1CheXbert using minimal additional computation over standard RAG.
4 Evaluation Strategy
To evaluate and compare LaB-RAG, we adopt RRG-specific metrics, F1-RadGraph [27] and F1-CheXbert [70]. These measure clinical relevance by computing the F1-score between clinical annotations of the actual and generated reports. The CheXbert annotations are the binarization of the 14 CheXpert [25] labels, while RadGraph annotations are broader in scope. We use the radgraph v0.1.12 and f1chexbert v0.0.2 python packages for their implementations of RadGraph and CheXbert. We also measure natural language metrics, using huggingface evaluate v0.4.2 to compute BLEU-4 [60], ROUGE-L [43], and BERTScore [82]. We refer the reader to our code and Sec. S4 for our specific usage of these metrics.
4.1 Comparison to Literature Models
We compare against both retrieval-based (CXR-RePaiR [18], CXR-ReDonE [64], X-REM [28]) and direct latent image embedding tuned models (CXRMate [56, 57], CheXagent [12], RGRG [73]). For CXRMate, we specifically use the version submitted to the BioNLP workshop [15, 57]. As each method may have slightly different filtering strategies for data preprocessing, we evaluate on the intersection of the test data splits for each method. LaB-RAG’s data preprocessing only requires there be a ground truth reference text to compare against, thus our test split is a superset of other methods’ test splits. More detailed descriptions of preparing and running each method from the literature are presented in Sec. S6.
As discussed in Sec. 2, RRG models can be categorized by the report section they generate. LaB-RAG is able to generate any target section given a corresponding retrieval corpus and thus comparison to other models only depend on code and data availability. We report results on generated “Findings”, “Impression”, and the concatenation of these sections generated independently. All comparison models were trained in-part over MIMIC-CXR [35].
FactMM-RAG [72] and CXR-RePaiR-Gen [65] are also retrieval based methods, however they do not share code. While BioViL-T [2] and BioViL [4] report RRG metrics and are usable as encoders, they do not share code for autoregressive or precise retrieval based generation. Finally, the only missing component of Pragmatic Retrieval/Llama [54] is code to extract the required indication section. As the MIMIC Code Repository [34] requires modification to extract indications, we are unable to exactly replicate Pragmatic’s precise method.
5 Studies and Experiments on LaB-RAG
In the following sections, we present results and analyses of our studies on LaB-RAG. These include comparisons of LaB-RAG to literature models and experiments on varying components of our framework. Tab. S10 is a full table of our results across all metrics, models, and experiments. We discuss limitations and next steps where relevant.
5.1 Baseline Comparisons
We consider models from the literature as our baseline. For models which we could compare against that generate the “Findings” section, none are retrieval based. On “Findings” generation (Fig. 3 Top), LaB-RAG does comparable to SFT methods (RGRG [73], CheXagent [12], CXRMate [57]) on F1CheXbert, however underperforms in F1RadGraph. Yet on “Impression” generation (Fig. 3 Bottom), LaB-RAG is not significantly worse than any SFT method (CheXagent, CXRMate) and outperforms previous retrieval-based methods (CXR-RePaiR [18], CXR-ReDonE [64], X-REM [28]). Across natural language metrics (Fig. S3), differences in performance, while statistically significant, are small in magnitude (Tab. S10).
Actual AP view of the chest. Right PICC (supp. dev.) is seen with tip at the upper SVC. Relatively low lung volumes are seen. The lungs however remain clear without consolidation, effusion or pulmonary vascular congestion. Cardiac silhouette appears moderately enlarged (cardiomegaly), likely accentuated due to low lung volumes and AP technique.
LaB-RAG In comparison with the previous radiograph, there is little change in the appearance of the heart and lungs. Cardiac silhouette is mildly prominent (cardiomegaly) and there is dense calcification within the aortic arch and descending portion. Blunting of the costophrenic angle is again seen, but no evidence of acute focal pneumonia or vascular congestion. Multiple triple-lumen catheter (supp. dev.) extends to the right atrium.
CXRMate Single portable view of the chest was compared to previous exam from ___. Right-sided PICC (supp. dev.) is seen with tip in the upper SVC. The lungs are clear of focal consolidation or pulmonary vascular congestion. The cardiac silhouette is enlarged (cardiomegaly) but stable in configuration. There is no large pleural effusion. There is no pneumothorax. Osseous and soft tissue structures are unremarkable.
Table 2: Example “Findings” section demonstrating overvaluation by F1CheXbert. Blue entities are annotated with their corresponding CheXbert label and are also extracted by RadGraph. Red entities are only extracted by RadGraph. Despite CXRMate attaining higher F1RadGraph than LaB-RAG ($0.6$ vs $0.2$) on this example, LaB-RAG matches CXRMate in F1CheXbert (both $1.0$) due to LaB-RAG’s usage of labels.
5.2 Variations of LaB-RAG
We experiment with the core label-boosting components of LaB-RAG and other modular components. Full results across all metrics are available in Sec. S7.
Label Filter and Format Effects. We first consider the simplest strategies for using labels, either “Exact” label match filtering or “Simple” positive label formatting. We show that inclusion of labels marginally changes F1RadGraph results while improving F1CheXbert performance by ${\sim}5\%$ over standard RAG for both “Findings” and “Impression” (Fig. 4). A comparable gain in performance from the literature between CheXagent [12] and CXRMate [57] on “Findings” generation (Fig. 3 Top) required complex new training strategies and additional datasets. As previously observed, LaB-RAG is also comparable to CXRMate in absolute measures. For LaB-RAG, we attain this improvement over standard RAG with minimal additional computation, only requiring “Simple” label formatting; however we do not observe a meaningful improvement when adding “Exact” label filtering to “Simple” formatting, the default LaB-RAG settings. We hypothesize that this is because the LLM is able to infer which parts of the example reports are useful to describe those labels.
To test this, we present inexact label matched examples to the LLM by using our “Partial” label filter (Sec. 3.4.1). Fig. S5 shows that with the default “Simple” prompt, the filters are not meaningfully different in performance. We confirm that inexactly labeled examples are selected by examining the image similarity rank of the filtered selections. In Fig. S10, we observe the “Exact” filter selects fewer than the “Partial” filter of the most visually similar examples. Thus the “Partial” filter will present more mismatched labels to the LLM, yet the stable performance implies the language model is attending to only relevant parts of the examples. Ultimately, given the “Exact” label filter’s improvement over standard RAG and the lack of detriment for the “Simple” label format, we use both label boosting components as defaults for LaB-RAG, aiming to potentially make it simpler for the LLM to parse high-quality label-matched text in the prompt.
Related to prompt complexity, we find that increasing the verbosity of the label format beyond positive labels with the “Simple” prompt tends to decrease performance (Fig. S6).
Image and Language Model Choice. As LaB-RAG is a fully modular framework, we analyze the choice of our pretrained models. We first compare comparable Mistral 7B LLMs: Instruct-v0.1 [29], BioMistral (Instruct-v0.1 further pretrained on PubMed Central) [41], and Instruct-v0.3 [50]. Fig. S7 shows that newer model versions improve RRG performance, while biomedical domain adaptation may be detrimental. This is contrary to previous literature [21] on biomedical domain adaptation, leading us to hypothesize that for a language-intensive approach (ICL), natural language capabilities may be more important than domain specific knowledge. One next step would be to test an LLM with domain-specific instruction following capabilities.
Additionally, the choice of image model directly impacts our label quality. We simulate the effect of using higher accuracy labels by using ground truth labels. We show in Fig. S8 that using true labels improves performance across the board, particularly in impression generation. The gap between LaB-Classifier predicted and true labels (Fig. S2) may be explained by Fig. S10, as image embedding similarity does not necessarily align with the CheXpert [25] labels. A natural extension may be to repeat our experiments with different pretrained image encoders or label sets.
Joint Section Evaluation. Finally, we experiment with modular retrieval examples. Given concatenations of the “Findings” and “Impression” for retrieval, LaB-RAG can jointly generate both sections. We compare this against performance of generating sections independently for studies which have both sections (Fig. S9 Top), showing that LaB-RAG does marginally better when generating jointly. We also compare against CheXagent [12] and CXRMate [57] by evaluating the concatenation of sections each method generates independently (Fig. S9 Bottom). We find a similar result to Fig. S3, where LaB-RAG is comparable to these methods in F1CheXbert but not on other metrics.
5.3 Extended Analyses
Motivated by the results comparing LaB-RAG to literature (Fig. 3), we sought to analyze why we were comparable to SOTA in F1RadGraph for “Impression” generation but not “Findings” generation. Tab. 2 includes an example of an actual “Findings” section and its corresponding generations by LaB-RAG and CXRMate [57]. We annotate the entities identified by CheXbert [70] with their label. These entities are also identified by RadGraph [27], which also identifies many more. Examining RadGraph annotations, we observe that the entities of the actual report more closely align with those of CXRMate’s generation, hence the F1RadGraph achieved for CXRMate was $0.6$ vs Lab-RAG’s $0.2$. However, we also observe that the computed CheXbert labels of both LaB-RAG and CXRMate exactly match those of the actual report, resulting in F1CheXbert of $1.0$.
We argue that F1CheXbert overvalues these generated “Findings” because it relies solely on 14 labels that describe disease related findings. This is most relevant to novel composite RRG metrics, such as RadCliQ [79], which include CheXbert as a component metric. As LaB-RAG explicitly uses these same labels in its label filter and format, we effectively ensure that these labels will be present in the generated report. This may also be the reason for LaB-RAG’s improved F1RadGraph performance in “Impression” generation. Since the “Impression” describes summary diagnoses, CheXbert and RadGraph entities may be more closely aligned with each other. It is thus an open question of which label set to use for label boosting. An interesting possibility may be using RadGraph derived labels.
5.4 Other Limitations
Though we make no claim of LaB-RAG’s clinical viability, an analysis of hallucinations is an important step towards this discussion for any model; Tab. 2 certainly shows hallucinations in both generations. Another potential limitation arises from the CheXpert [25] labels included by MIMIC-CXR-JPG [32]; these labels were derived on mixed report sections, however we do not consider their source for our section-specific generation. Lastly, a limitation of all methods using pretrained LLMs is the difficulty of detecting data leakage from pretraining over unknown training data.
6 Conclusion
In this work, we present LaB-RAG: a modular framework for image captioning using pretrained models. We study and analyze LaB-RAG in the context of RRG, showing that it achieves SOTA on clinical language metrics. We offer insights into the importance of different components of LaB-RAG, while also critiquing a widely used RRG metric, arguing we can inflate its results. Throughout our experiments, we stress the simplicity of LaB-RAG as an ensemble of frozen DL models boosted by classical linear models. The key to LaB-RAG is leveraging categorical labels as natural descriptors of images. Doing so enables LaB-RAG to use off-the-shelf models which do not need to be jointly trained, a task which requires large paired datasets. This enables LaB-RAG to be fully modular and flexible for users.
Supplementary Material
S1 Abbreviations and Terms
Term Meaning
LaB-RAG Label Boosted Retrieval Augmented Generation
Label Filtering Using categorical labels to filter retrieved examples
Label Formatting Using categorical labels as text descriptors of images
RRG Radiology report generation
CXR Chest X-ray
AP Anterior-posterior (i.e. from front to back)
PICC Peripherally inserted central catheter
SVC Superior vena cava
AI Artificial intelligence
ML Machine learning
DL Deep learning
LLM Large language model
RAG Retrieval augmented generation
FM Foundation model
SFT Supervised fine-tuning
PEFT Parameter-efficient fine-tuning
ICL In-context learning
SOTA State of the art
NLP Natural language processing
CLIP Contrastive language-image pretraining
Table S1: Reference of abbreviations.
S2 Dataset Extended Information
MIMIC-CXR [35] imaging studies were collected in the emergency department at the Beth Israel Deaconess Medical Center (BIDMC) in Boston, MA between 2011 and 2016. Multiple chest X-rays may be present for a single study, and a patient may have multiple imaging studies. The training, validation, and testing splits provided by MIMIC-CXR are split at the patient level, meaning all images for all studies belonging to a single patient are contained in one split. The demographic breakdown of studies in each split is presented in Tab. S2, though the final number of patients and imaging studies considered in each of our experiments depends on the availability of all required data modalities. Demographic information was collected from the MIMIC-IV v2.2 [37, 31] hosp module and MIMIC-IV-ED v2.2 [30]. All data were accessed through PhysioNet [20].
For images, we select one image per study by preferentially selecting images based on the captured view position. The number of selected image views per split is listed in Tab. S3. We follow the order of preference for view positions from the ACL 2024 BioNLP workshop’s RRG shared task [15]; the exact list can be found in Tab. S4. We use code inspired by the MIMIC Code Repository [34] to extract radiology report sections from the free text reports. The number of extracted findings and impression sections per split is listed in Tab. S5.
Ethical Considerations for Data Usage. As the MIMIC [37] datasets were collected from the BIDMC, the BIDMC institutional review board (IRB) granted a waiver of informed consent and approved the collection and sharing of the data for research purposes following proper de-identification according to HIPAA. As users of the data, we follow procedures outlined by PhysioNet [20] to have all researchers with data access complete human research training and sign a data use agreement (DUA). We strictly control access to the data in our compute environments and make no efforts to re-identify data subjects from any free-text data (which may have escaped de-identification) or any other sources. Additionally, as we do no fine-tuning of our own LLMs, we do not need to consider whether it is a DUA violation to share LLM weights which may have memorized and can reproduce training text.
S3 LaB-Classifier Training Details
We fit 14 per-label binary logistic regressions over the unprojected 512d image embeddings from BioViL-T (see Sec. 3.2). We derive ground-truth labels to train our models with by binarizing the CheXpert labels as positive (value of 1) or not (value of 0, -1, or null). We fit the models on the training split and find a per-label probability threshold which maximizes the f1-score over the validation split. The receiver operator characteristic and precision recall curves of the 14 labels over the test split are presented in Fig. S2. We adopt the same heuristic as above for computing the “Other” label for the image. For training the logistic classifiers, we use sklearn [61] v1.5.1 with L2 regularization, the saga solver for 500 iterations, and otherwise default hyperparameters.
Figure S1: Overview of training per-label LaB-Classifier logistic regressions.
Figure S2: Test set ROC and PR curves of LaB-Classifier, a set of 14 per-label logistic regressions.
S4 Evaluation Strategy Extended Details
For all metrics, actual and generated text are input as whole reports; reports are not split on the sentence level. This gives a score between 0 and 1 for each report for each metric. We compare results by considering the per-metric scores across all reports. We show barplots of the per-metric average scores with errorbars showing the standard error. To test for statistical significance, we perform paired t-tests within one set of experiments, applying Bonferroni correction for the number of comparisons made within a single metric. For experiments on variations of LaB-RAG, we compare all pairwise combinations of the variants; for comparing to literature models, we only consider pairs including LaB-RAG. Statistically significant pairs are annotated with brackets in the barplot with ‘*’ denoting $p<0.05$; nonsignificant pairs are not annotated. We refer to our code for precise implementation details.
Overall Train Val Test
Count, N Patients 65379 64586 500 293
Studies 227835 222758 1808 3269
Age Median [Q1,Q3] 65 [52,76] 65 [52,76] 63 [52,77] 69 [61,77]
Missing, N (%) 11825 (5.2) 11522 (5.2) 134 (7.4) 169 (5.2)
Sex, N (%) Female 105117 (48.7) 102951 (48.7) 737 (44.0) 1429 (46.1)
Male 110893 (51.3) 108285 (51.3) 937 (56.0) 1671 (53.9)
Missing 11825 (5.2) 11522 (5.2) 134 (7.4) 169 (5.2)
Race, N (%) White 141531 (65.5) 138343 (65.5) 1081 (64.6) 2107 (68.0)
Black 35788 (16.6) 34920 (16.5) 186 (11.1) 682 (22.0)
Hispanic or Latino 12358 (5.7) 12150 (5.8) 110 (6.6) 98 (3.2)
Other 9088 (4.2) 8857 (4.2) 190 (11.4) 41 (1.3)
Unknown 8430 (3.9) 8343 (4.0) 51 (3.0) 36 (1.2)
Asian 7927 (3.7) 7753 (3.7) 47 (2.8) 127 (4.1)
AIAN 644 (0.3) 628 (0.3) 7 (0.4) 9 (0.3)
NHPI 205 (0.1) 203 (0.1) 2 (0.1) 0 (0.0)
Missing 11864 (5.2) 11561 (5.2) 134 (7.4) 169 (5.2)
Table S2: Per-study demographic breakdown of MIMIC defined training/validation/testing splits. AIAN = American Indian or Alaska Native, NHPI = Native Hawaiian or Pacific Islander.
View, N (%) Overall Train Val Test
PA 85867 (37.7) 84323 (37.9) 661 (36.6) 883 (27.0)
AP 132264 (58.1) 129034 (57.9) 1072 (59.3) 2158 (66.0)
LATERAL 176 (0.1) 172 (0.1) 2 (0.1) 2 (0.1)
LL 2225 (1.0) 2160 (1.0) 17 (0.9) 48 (1.5)
AP AXIAL 1 (0.0) 1 (0.0) 0 (0.0) 0 (0.0)
LAO 1 (0.0) 1 (0.0) 0 (0.0) 0 (0.0)
LPO 1 (0.0) 1 (0.0) 0 (0.0) 0 (0.0)
Table S3: Count of selected single view per-study. PA = postero-anterior, AP = antero-posterior, LL = left lateral, LAO = left anterior oblique, LPO = left posterior oblique.
Rank View Rank View Rank View
1 PA 6 AP LLD 11 RAO
2 AP 7 AP RLD 12 LPO
3 LATERAL 8 PA RLD 13 XTABLE LATERAL
4 LL 9 PA LLD 14 SWIMMERS
5 AP AXIAL 10 LAO 15 (Empty String)
Table S4: Order of view preference for selecting single image for multi-image studies. PA = postero-anterior, AP = antero-posterior, LL = left lateral, LLD = left lateral decubitus, RLD = right lateral decubitus, LAO = left anterior oblique, RAO = , LPO = left posterior oblique.
Section, N (%) Overall Train Val Test
Impression 189561 (83.2) 185816 (83.4) 1521 (84.1) 2224 (68.0)
Findings 155716 (68.3) 152173 (68.3) 1196 (66.2) 2347 (71.8)
Both 128032 (56.2) 125417 (56.3) 991 (54.8) 1624 (49.7)
Table S5: Count of reports with extracted “Findings”, “Impression”, or both sections.
S5 Prompts
Prompt Type Prompt Text
Naive You are an expert radiological assistant. Your task is to generate a radiology report after <<Report>> given context information. The context information contains examples of reports written for similar cases. Use the examples to generate a report for the current case. Strictly follow the instructions below to generate the reports. *Instructions** 1. The report must be based on the information in the context. 2. The report must mimic the style of the reports shown in the context. 3. Do not generate blank reports. CONTEXT: Example: 1 (Report Text) (More Examples) Now generate the report for the current case. Always generate reports based on the examples shown. <<Report>>
Table S6: The “Naive” prompt does not incorporate labels.
Prompt Type Prompt Text
Simple You are an expert radiological assistant. Your task is to generate a radiology report after <<Report>> given context information. The context information contains examples of reports written for similar cases and their associated labels. Use the examples and their associated labels to generate a report for the current case based on the current label. Strictly follow the instructions below to generate the reports. *Instructions** 1. The report must be based on the information in the context and the current label. 2. The report must mimic the style of the reports shown in the context. 3. Do not generate blank reports. CONTEXT: Example: 1 Label: (Positive Labels) (Report Text) (More Examples) Now generate the report for the current case using its label below. Always generate reports based on the examples shown. Label: (Positive Labels) <<Report>>
Table S7: The “Simple” prompt formats positive labels for the examples and target image.
Prompt Type Prompt Text
Verbose You are an expert radiological assistant. Your task is to generate a radiology report after <<Report>> given context information. The context information contains examples of reports written for similar cases and their associated labels. The labels provided are expert annotations. More information about the labels is described below. The individual labels used represent common chest radiographic observations and fall under four categories: ‘Positive’, ‘Negative’, ‘Uncertain’ and ‘Unmentioned’. These categories correspond to the mention or presence of the labels or their equivalent in the report. Below is a description and example of each of the label categories: 1. ‘Positive’: A label is positive if the associated observation or disease is stated as present        in the report, for example: ‘moderate bilateral effusions observed’. 2. ‘Negative’: A label is negative if the associated observation or disease is stated as absent        in the report, for example: ‘no evidence of pulmonary edema’. 3. ‘Uncertain’: A label is uncertain if there is ambiguity about the presence or absence of        the associated observation or disease in the report, for example: ‘pneumonia cannot be excluded        in the appropriate clinical context’. 4. ‘Unmentioned’: A label is unmentioned if there is no mention of the associated observation        or disease in report. Use the examples, their associated labels, and the label descriptions to generate a report for the current case based on the current label. Strictly follow the instructions below to generate the reports. *Instructions** 1. The report must be based on the information in the context and the current label. 2. The report must mimic the style of the reports shown in the context. 3. Do not generate blank reports. CONTEXT: Example: 1 Positive: (Positive Labels) Negative: (Negative Labels) Uncertain: (Uncertain Labels) Unmentioned: (Unmentioned Labels) (Report Text) (More Examples) Now generate the report for the current case using its label below. Always generate reports based on the examples shown. Positive: (Positive Labels) Negative: (Negative Labels) Uncertain: (Uncertain Labels) Unmentioned: (Unmentioned Labels) <<Report>>
Table S8: The “Verbose” prompt formats positive, negative, uncertain, and unmentioned labels for the examples and target image.
Prompt Type Prompt Text
Instruct (Same as Verbose) *Instructions** 1. The report must be based on the information in the context and the current label. 2. The report must mimic the style of the reports shown in the context. 3. Do not generate blank reports. 4. Ensure that the positive labels are clearly described as being present in the report,        using example language from the context. 5. Ensure that the negative labels are clearly described as being absent in the report,        using example language from the context. 6. Describe the uncertain labels as necessary. 7. Ensure that the unmentioned labels are not mentioned in the report. (Same as Verbose)
Table S9: The “Instruct” prompt uses the same label format as “Verbose” and adds additional instructions.
S6 Literature Model Inference
In this section, we provide our extended descriptions for baseline models from the literature. We select models based on model architecture, generated report section, SOTA performance on evaluation metrics, and finally availability of open source inference code. The final set of models we compare against are as follows: CXR-RePaiR [18], CXR-ReDonE [64], X-REM [28], RGRG [73], CheXagent [12], and CXRMate-RRG24 [57]. We compare models over the intersection of each method’s test split subsets.
CXR-RePaiR formulates report generation as a pure retrieval task. It uses a model, trained via CLIP on radiology report-image pairs from MIMIC-CXR, to rank similarity between test images and a large retrieval corpus. We use the mode of CXR-RePaiR where the generated report output is the the top-1 retrieved whole reports.
CXR-RePaiR generates the “Impression” section over the MIMIC-CXR test split. Only the subset of the test split with an extractable “Impression” section is considered. While we adopt a similar strategy, CXR-RePaiR uses a custom implementation of “Impression” section extraction that differs from ours. CXR-RePaiR thus uses 2192 test studies as compared to our “Impression” test split of 2224.
CXR-ReDonE improves upon CXR-RePaiR by preprocessing the training and retrieval reports to remove references to prior imaging studies. Additionally, the joint-embedding model was trained via ALBEF instead of CLIP. We use the provided checkpoint of the retrieval model which was trained over the new data and via the new method; we do inference using whole reports with priors omitted as the retrieval set. The generated reports are then the top-1 retrieved whole report. CXR-ReDonE uses the “Impression” sections preprocessed by CXR-RePaiR and so results in the same test split subset.
X-REM also trains a vision-language model via ALBEF, however they introduce a novel similarity metric during training which incorporates CheXbert labels. An intermediary step retrieves the top-10 whole reports ranked by their new similarity score. Finally, they apply a model-based natural language filter to each retrieved report and only select those that are deemed relevant up to a limit; we adopt the author default limit of 2 reports, thus the output report is a concatenation of up to 2 reports. X-REM also uses the “Impression” sections preprocessed by CXR-RePaiR with the same test split subset.
CheXagent is presented by its authors as a foundation model trained to follow instructions in the CXR domain. CheXagent follows a complex training scheme and utilizes other CXR datasets. At a high-level, CheXagent is trained by aligning image latent embeddings to the an LLM’s token embedding space. Additionally, the authors introduce a novel dataset for instruction tuning, their final training step.
CheXagent is able to generate both “Findings” and “Impression” sections. “Impression” generated is simply prompted with “Generate impression”. “Findings” are generated by concatenating generations of individual “Local Findings” per anatomical compartment. “Local Findings” are generated by prompting with “Describe [Anatomical Compartment]” where the compartments are: “Airway”, “Breathing”, “Cardiac”, “Diaphragm” and “Everything else (e.g., mediastinal contours, bones, soft tissues, tubes, valves, and pacemakers)”. These compartments were inspired by documentation provided by the authors. CheXagent inference code is flexible and allows for generation over our specified subsets of the test split for “Findings”, “Impression”, or both sections.
CXRMate uses an encoder-decoder transformer architecture. It is both able to generate a single report over multiple images from a single imaging study and it takes as input reports from prior studies. Additionally, it is trained with a complex reinforcement learning framework. Interestingly, the authors experiment with RadGraph and CXR-BERT in their reward function, arguing that CXR-BERT better captures radiology report semantics; CXR-BERT is the language encoder of the vision-langauge model BioViL [4].
We specifically use the checkpoint of CXRMate submitted to the ACL 2024 BioNLP workshop’s task for RRG [57], though we only input the single image per study defined by our data preprocessing steps. While CheXagent generates both “Findings” and “Impression” jointly, the model provides a utility to split these sections after generation. We are then able to run inference using our subsets of the test split for which we have “Findings”, “Impression”, or both sections.
RGRG generates the “Findings” section of a report by combining individual sentences that describe specific anatomical regions. This is accomplished by training a model over a specialized dataset, derived from MIMIC-CXR, which pairs anatomical region bounding boxes with sentences from the corresponding report detailing those regions. Thus RGRG learns to extract localized image latent embeddings to generate sentences grounded in the specific anatomical feature. As RGRG’s language model is a decoder-only transformer, the image embeddings are prepended input as tokens. Unlike CheXagent, where individual anatomical regions must be prompted by the user, RGRG automatically selects relevant regions using a custom trained object detector model. As RGRG only generates the “Findings”, we evaluate using our test split subset for studies with an extractable “Findings” section.
S7 Supplemental Results
Experiment Section Variable bleu4 rougeL bertscore f1radgraph f1chexbert
Baselines Findings LaB-RAG 0.042 0.204 0.861 0.183 0.466
RGRG 0.071 0.210 0.860 0.215 0.435
CheXagent 0.061 0.232 0.861 0.234 0.412
CXRMate 0.064 0.229 0.861 0.247 0.469
Impression LaB-RAG 0.042 0.204 0.861 0.183 0.466
CXR-RePaiR 0.020 0.124 0.848 0.108 0.347
CXR-ReDonE 0.009 0.130 0.845 0.103 0.333
X-REM 0.013 0.162 0.854 0.139 0.408
CheXagent 0.060 0.209 0.864 0.176 0.397
CXRMate 0.053 0.212 0.862 0.191 0.422
Both LaB-RAG 0.042 0.204 0.861 0.183 0.466
CheXagent 0.093 0.263 0.870 0.272 0.415
CXRMate 0.087 0.253 0.866 0.271 0.438
Filter & Format Findings Standard RAG 0.043 0.201 0.858 0.183 0.409
Label Filter only 0.041 0.198 0.857 0.180 0.443
Label Format only 0.045 0.208 0.863 0.193 0.466
LaB-RAG 0.042 0.204 0.861 0.183 0.466
Impression Standard RAG 0.020 0.149 0.851 0.137 0.382
Label Filter only 0.023 0.152 0.851 0.144 0.438
Label Format only 0.033 0.172 0.858 0.152 0.446
LaB-RAG 0.032 0.168 0.856 0.146 0.450
Filter Findings No-filter 0.045 0.208 0.863 0.193 0.466
Exact 0.042 0.204 0.861 0.183 0.466
Partial 0.043 0.205 0.862 0.185 0.467
Impression No-filter 0.033 0.172 0.858 0.152 0.446
Exact 0.032 0.168 0.856 0.146 0.450
Partial 0.033 0.174 0.858 0.154 0.451
Format Findings Naive 0.041 0.198 0.857 0.180 0.443
Simple 0.042 0.204 0.861 0.183 0.466
Verbose 0.040 0.196 0.856 0.180 0.438
Instruct 0.039 0.195 0.855 0.180 0.433
Impression Naive 0.023 0.152 0.851 0.144 0.438
Simple 0.032 0.168 0.856 0.146 0.450
Verbose 0.013 0.147 0.847 0.134 0.400
Instruct 0.013 0.146 0.846 0.131 0.393
LLM Findings Mistral-v1 0.037 0.178 0.850 0.181 0.454
BioMistral 0.038 0.197 0.857 0.168 0.445
Mistral-v3 0.042 0.204 0.861 0.183 0.466
Impression Mistral-v1 0.017 0.124 0.838 0.139 0.446
BioMistral 0.029 0.160 0.851 0.131 0.402
Mistral-v3 0.032 0.168 0.856 0.146 0.450
Label Findings True 0.055 0.221 0.866 0.203 0.568
Predicted 0.042 0.204 0.861 0.183 0.466
Impression True 0.047 0.229 0.871 0.221 0.800
Predicted 0.032 0.168 0.856 0.146 0.450
Section Intersection Findings-Intersect 0.050 0.214 0.863 0.197 0.431
Impression-Intersect 0.039 0.182 0.860 0.151 0.430
Both 0.061 0.226 0.865 0.211 0.456
Table S10: Full experimental results. By default, LaB-RAG uses “Exact” filter, “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S3: Comparison of LaB-RAG to literature baselines. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of comparisons vs LaB-RAG. By default, LaB-RAG uses “Exact” filter, “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S4: Comparison of label filter and format components of LaB-RAG. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. “Standard RAG” uses “No-filter” filter and “Naive” format/prompt. “Label Filter only” uses “Exact” filter and “Naive” format/prompt. “Label Format only” uses “No-filter” filter and “Simple” format/prompt. LaB-RAG otherwise defaults to “Exact” filter, “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S5: Comparison of LaB-RAG label filters. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S6: Comparison of LaB-RAG label formats. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Exact” filter, Mistral-v3, and predicted labels.
Figure S7: Comparison of different LLM modules for LaB-RAG. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Exact” label filter, “Simple” format/prompt, and predicted labels.
Figure S8: Comparison of label quality’s effect on LaB-RAG. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Exact” format, “Simple” format/prompt, and Mistral-v3.
Figure S9: Comparison of generating different report sections and comparison of joint section generation against literature models. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Exact” format, “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S10: Histogram of image similarity rank for top 5 studies selected for each inference image, after “Exact” label filtering or “Partial” label reordering. While similar images may share some labels with the inference image, they may not match exactly.